OpenStudy (alina123):

find the number of full 8 hour shifts that jenna worked last month. she worked twice as many 6 hour shifts as 8 hour shifts she worked a total of 280 hour last month

OpenStudy (zzr0ck3r):

are we assuming that there are only 8 hour and 6 hour shifts?

OpenStudy (alina123):

I guess yes

OpenStudy (mathstudent55):

Have you learned systems of equations?

OpenStudy (alina123):

No

OpenStudy (mathstudent55):

Ok. Since we don't know the number of 6-hour or 8-hour shifts, we need a variable to represent the number of 8-hour shifts. Let's call x the number of 8-hourt shifts.

OpenStudy (mathstudent55):

If she worked twice as many 6-hour shifts as 8-hour shifts, and she worked x 8-hour shifts, then using x, how many 6-hour shifts did she work?

OpenStudy (alina123):

2

OpenStudy (mathstudent55):

Close. Not just 2, but 2 times x or 2x.

OpenStudy (mathstudent55):

Ok?

OpenStudy (alina123):

Yup

OpenStudy (mathstudent55):

Ok. We need one more step. then we can write an equation. In an 8-hour shift, she works 8 hours. This is obvious. In 2 8-hour shifts, she works 2 * 8 hours = 16 hours. In 3 8-hour shifts, she works 3 * 8 hours = 24 hours. In x 8-hour shifts, she works how many hours?

OpenStudy (alina123):

8x or 8 hours

OpenStudy (mathstudent55):

No. You got it right the first time. In x 8-hour shifts, she works x * 8 hours, or 8x hours.

OpenStudy (alina123):

Ok so 8 times x is 8x

OpenStudy (mathstudent55):

Now we can put all the information together and write an equation.

OpenStudy (mathstudent55):

Correct.

OpenStudy (mathstudent55):

She worked x 8-hour shifts. That means she worked 8x hours in 8-hour shifts. She worked 2x 6-hour shifts. That means she worked 2x * 6 = 12x hours in 6-hour shifts. She worked a total of 280 hours. That means that 8x + 12x = 280

OpenStudy (mathstudent55):

Now we solve that equation to find x, the number of 8-hour shifts she worked.

OpenStudy (mathstudent55):

Since 8x and 12x are like terms, you can add them together.

OpenStudy (alina123):

to get 20x

OpenStudy (mathstudent55):

Correct. Now we have 20x = 280 We divide both sides by 20 to find x.

OpenStudy (alina123):

and you get 14 which would represent 14, 8 hour shifts

OpenStudy (mathstudent55):

Correct.

OpenStudy (mathstudent55):

Now let's see if the answer makes sense.

OpenStudy (mathstudent55):

Remember she worked twice as many 6-hour shifts as she worked 8-hour shifts. That means she worked 14 8-hour shifts and 28 6-hour shifts, right?

OpenStudy (alina123):

how did you get 28 6 hour shifts

OpenStudy (alina123):

nevermind i see 14 times 2

OpenStudy (mathstudent55):

Good.

OpenStudy (alina123):

because its twice the time

OpenStudy (mathstudent55):

Now let's see if the total number of hours really adds up to 280. 14 * 8 = 112 28 * 6 = 168 112 + 168 = 280 Yes, the number of hours adds up to 280, so our answer is correct.

OpenStudy (mathstudent55):

Ok, gtg. Bye.

OpenStudy (alina123):

ok thanks i first thought that since twice 6 hour shifts then its 12 just divided by 280 but you have to include the 8 to

OpenStudy (mathstudent55):

Correct. You need to account for both the 6-hour and 8-hour shifts.

OpenStudy (alina123):

yup ok